OpenStudy (anonymous):

the function g is given by g(x)=3x^2/e^(3x). On which of the following intervals is g increasing? a)(-∞,0) b)(-∞,2/3) c)(0,2/3) d)(0,∞) e)(2/3,∞)

OpenStudy (anonymous):

did you find the derivative?

OpenStudy (anonymous):

no I'm not sure how to find the derivative of the denominator..

OpenStudy (danjs):

Here is the graph for reference before doing the work to find critical points...

OpenStudy (danjs):

You can use the product rule on this one, rewrite the e^(3x) term in the numerator first. \[g(x) = 3x^2*e^{-3x}\]

OpenStudy (danjs):

The product rule gives. \[g '(x) = \frac{ d }{ dx }g(x) = 3x^2*\frac{ d }{ dx }e^{-3x}~~+~~e^{-3x}*3\frac{ d }{ dx }x^2\]

OpenStudy (danjs):

Can you do those individual derivatives?

OpenStudy (anonymous):

not to be argumentative, but if you think you will have an easier time with the quotient rule, you will have to pay for it in the end when you add up the fractions, which you must do to answer the question

OpenStudy (anonymous):

idk wat the derivative of e^3x is

OpenStudy (danjs):

I changed it to use the product rule, quotient?

OpenStudy (anonymous):

sorry i meant you need to use the quotient rule the derivative of \(e^{3x}\) is \[3e^{3x}\] by the chain rule

OpenStudy (anonymous):

you gotta get the numerator no matter what rule you use, so the product rule is a bad shortcut here

OpenStudy (anonymous):

okay so g'(x)=e^3x(6x)-3x^2(3e^3x)/(e^3x)^2

OpenStudy (anonymous):

ignore the denominator, factor the numerator

OpenStudy (danjs):

To each their own. It was pretty easy this way too i think, after factoring. \[g'(x) = 3xe^{-3x}[2-3x] =0\]

OpenStudy (anonymous):

then drop the \(e^{-3x}\) since that is always positive, and concentrate on \[x(2-3x)\]

OpenStudy (anonymous):

okay..then what?

OpenStudy (anonymous):

can you figure out where \[x(2-3x)>0\) ?

OpenStudy (anonymous):

oops \[x(2-3x)>0\]

OpenStudy (anonymous):

uhm -1

OpenStudy (anonymous):

hmm your answer needs to be an interval hint, \[y=x(2-3x)\] is parabola that opens down it is positive between the zeros

OpenStudy (anonymous):

uhm well wouldn't it be greater than zero at any negative number?

OpenStudy (anonymous):

where is it zero?

OpenStudy (danjs):

Set the derivative = 0. This will tell you possible critical points to check for. \[g'(x) = 3xe^{-3x}[2-3x]=0\] This is true if: \[3xe^{-3x}=0~~~~~or~~~~~2-3x=0\] \[x=0~~~~~or~~~~~x=2/3\] set up a number line, and check the intervals for inc/dec

OpenStudy (danjs):



OpenStudy (danjs):

Pick a number in the intervals: (-infinity, 0) , (0, 2/3), and (2/3 , +infinity) and test it in g '(x) \[g '(x) >0~~~~increasing \] \[g'(x)<0~~~~~decreasing\]

OpenStudy (danjs):

good?, you remember it sort of now?

OpenStudy (anonymous):

okay so (2/3, ∞ ) it is decreasing?

OpenStudy (danjs):

The derivative tells you what the slope of a tangent line is at a certain point on the curve. If the derivative is positive, the tangent has a positive slope, from the image, you can imagine the curve must be increasing.

OpenStudy (anonymous):

& yes i can understand it better now thanks

OpenStudy (danjs):

i didn't calculate anything, but the graph is linked in my first post.... It is decreasing on \[(2/3, \infty)\] but try a number in g '(x) that is less than 0, what do you get?

OpenStudy (anonymous):

when it's less than zero, i get a positive number..

OpenStudy (danjs):

hmm, g '(-1) is \[g'(-1) = 3*(-1)e^{-3*(-1)}*[2-3(-1)] = -3e^3 * (2+3)\]

OpenStudy (danjs):

negative times a positive ----g '(-1) is a negative number

OpenStudy (danjs):



OpenStudy (anonymous):

oh ok i didn't plug it into the entire derivative. I disregarded a part of it. but it makes sense now. Thank you so much.

OpenStudy (danjs):

you're welcome... here are some things i found that may help ...

OpenStudy (danjs):

it is all open free stuff from the web site linked in the things..

OpenStudy (anonymous):

okay thanks again. I appreciate it :)